# Insert-Only versus Insert-Delete in Dynamic Query Evaluation

### v3

@@ E.1.1. From the Maintenance of $Q$ to the Maintenance of $\overline{Q}$
The following proposition will be useful later for the proof of Theorem 3.4. It basically says that when we truncate a tuple from $\overline{R}([\tau^{\prime},\infty],\bm{t})$ to $\overline{R}([\tau^{\prime},\tau],\bm{t})$, the truncated tuple is still going to join with all other tuples that it used to join with before the truncation.
@@ E.2. Upper Bound for $\text{IVM}^{\pm}_{\delta}[Q]$
In the previous section, we proved Theorem 3.4 for the $\text{IVM}^{\pm}[Q]$ problem for a given query $Q$. In this section, we extend the proof and the corresponding algorithm from the previous section in order to prove Theorem 3.4 for the $\text{IVM}^{\pm}_{\delta}[Q]$ problem.
@@
We call the above modified version of Algorithm 2 the delta version of Algorithm 2. Before we prove that the delta version of Algorithm 2 proves Theorem 3.4 for $\text{IVM}^{\pm}_{\delta}[Q]$, we need some preliminaries.
@@
The following theorem implies Theorem 3.4 for $\text{IVM}^{\pm}_{\delta}[Q]$:
@@
The proof is very similar to the proof of Theorem 3.4 from Section E.1 for $\text{IVM}^{\pm}[Q]$. For every insert or delete $\delta_{\tau}\mathcal{D}=\{sR_{j}(\bm{t})\}$ where $s\in\{+,-\}$, we now have to insert an extra tuple $+\overline{R}_{j}([\tau,\tau],\bm{t})$ into $\overline{\mathcal{D}}$. This only increases the size of $\overline{\mathcal{D}}$ by a constant factor. Hence, it doesn’t affect the overall runtime.
@@
The upper bound in Theorem 3.4 is written in terms of the length $N$ of the update stream. In this section, we prove Corollary 3.5, which gives a stronger bound where $N$ is replaced by the current database size $|\mathcal{D}^{(\tau)}|$ at the $\tau$-th update. Note that $|\mathcal{D}^{(\tau)}|\leq N$ and $|\mathcal{D}^{(\tau)}|$ could be unboundedly smaller than $N$ (specifically when the stream contains many deletes before the timestamp $\tau$).
Corollary 3.5.
For any query $Q$, both $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau)}|^{\mathsf{w}(\widehat{Q})-1})$ amortized update time, where $\widehat{Q}$ is the multivariate extension of $Q$, and $|\mathcal{D}^{(\tau)}|$ is the current database size at update time $\tau$.
@@
The proof uses Theorem 3.4 as a black box. In particular, we will use the same maintenance algorithm from the proof of Theorem 3.4. However, every now and then, we will “reset” the algorithm by taking the current database $\mathcal{D}^{(\tau)}$, creating a new empty database from scratch, and inserting every tuple in $\mathcal{D}^{(\tau)}$ into the new database as a single-tuple update. After that, we process the next $\frac{1}{2}|\mathcal{D}^{(\tau)}|$ updates [^7] For simplicity, we assume that $|\mathcal{D}^{(\tau)}|$ is a positive even integer. from the original update sequence, and then we apply another reset. The total number of updates between the last two resets is thus $\frac{3}{2}|\mathcal{D}^{(\tau)}|$. By Theorem 3.4, these $\frac{3}{2}|\mathcal{D}^{(\tau)}|$ updates together take time $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau)}|^{\mathsf{w}(\widehat{Q})})$. Moreover, note that between the two resets, the database size cannot change by more than a factor of 2. This is because we only have $\frac{1}{2}|\mathcal{D}^{(\tau)}|$ updates from the original stream that are applied to $\mathcal{D}^{(\tau)}$. In the two extreme cases, all of these updates are inserts or all of them are deletes.
@@
Eq. (62) holds because between $\tau_{i}$ and $\tau_{i+1}$, we are making $\frac{1}{2}|\mathcal{D}^{(\tau_{i})}|$ updates to $\mathcal{D}^{(\tau_{i})}$, which in the two extreme cases are either all inserts or all deletes. As noted before, from the beginning of the $\tau_{i}$-reset until the beginning of the $\tau_{i+1}$-reset, we process $\frac{3}{2}|\mathcal{D}^{(\tau_{i})}|$ updates, which together take time $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau_{i})}|^{\mathsf{w}(\widehat{Q})})$, thanks to Theorem 3.4. To bound the total runtime, we sum up over $i\in[k-1]$: (We don’t actually need to perform the last reset at time $\tau_{k}=N$ since there are no subsequent updates anyhow.)
@@ E.4. Lower Bound for $\text{IVM}^{\pm}_{\delta}[Q]$
The goal of this section is to prove Theorem 3.6, which gives a lower bound on the $\text{IVM}^{\pm}_{\delta}$ update time of a query $Q$ in terms of a lower bound on the static evaluation time for any component $\widehat{Q}_{\bm{\sigma}}$ in its multivariate extension. To that end, we give a reduction from the static evaluation of $\widehat{Q}_{\bm{\sigma}}$ to the $\text{IVM}^{\pm}_{\delta}$ problem for $Q$. Example 7.2 introduces the main idea behind this reduction.

### v4

@@ E.1.1. From the Maintenance of $Q$ to the Maintenance of $\overline{Q}$
The following proposition will be useful later for the proof of Lemma 7.1. It basically says that when we truncate a tuple from $\overline{R}([\tau^{\prime},\infty],\bm{t})$ to $\overline{R}([\tau^{\prime},\tau],\bm{t})$, the truncated tuple is still going to join with all other tuples that it used to join with before the truncation.
@@ E.2. Upper Bound for $\text{IVM}^{\pm}_{\delta}[Q]$
In the previous section, we proved Lemma 7.1 for the $\text{IVM}^{\pm}[Q]$ problem for a given query $Q$. In this section, we extend the proof and the corresponding algorithm from the previous section in order to prove Lemma 7.1 for the $\text{IVM}^{\pm}_{\delta}[Q]$ problem.
@@
We call the above modified version of Algorithm 2 the delta version of Algorithm 2. Before we prove that the delta version of Algorithm 2 proves Lemma 7.1 for $\text{IVM}^{\pm}_{\delta}[Q]$, we need some preliminaries.
@@
The following theorem implies Lemma 7.1 for $\text{IVM}^{\pm}_{\delta}[Q]$:
@@
The proof is very similar to the proof of Lemma 7.1 from Section E.1 for $\text{IVM}^{\pm}[Q]$. For every insert or delete $\delta_{\tau}\mathcal{D}=\{sR_{j}(\bm{t})\}$ where $s\in\{+,-\}$, we now have to insert an extra tuple $+\overline{R}_{j}([\tau,\tau],\bm{t})$ into $\overline{\mathcal{D}}$. This only increases the size of $\overline{\mathcal{D}}$ by a constant factor. Hence, it doesn’t affect the overall runtime.
@@
The upper bound in Lemma 7.1 is written in terms of the length $N$ of the update stream. In this section, we show how to use Lemma 7.1 as a black box in order to prove Theorem 4.4, which gives a stronger bound where $N$ is replaced by the current database size $|\mathcal{D}^{(\tau)}|$ at the $\tau$-th update. Note that $|\mathcal{D}^{(\tau)}|\leq N$ and $|\mathcal{D}^{(\tau)}|$ could be unboundedly smaller than $N$ (specifically when the stream contains many deletes before the timestamp $\tau$).
Theorem 4.4.
For any query $Q$, both $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau)}|^{\mathsf{w}(\widehat{Q})-1})$ amortized update time and non-amortized constant enumeration delay, where $\widehat{Q}$ is the multivariate extension of $Q$, and $|\mathcal{D}^{(\tau)}|$ is the current database size at update time $\tau$.
@@
The proof uses Lemma 7.1 as a black box. In particular, we will use the same maintenance algorithm from the proof of Lemma 7.1. However, every now and then, we will “reset” the algorithm by taking the current database $\mathcal{D}^{(\tau)}$, creating a new empty database from scratch, and inserting every tuple in $\mathcal{D}^{(\tau)}$ into the new database as a single-tuple update. After that, we process the next $\frac{1}{2}|\mathcal{D}^{(\tau)}|$ updates [^8] For simplicity, we assume that $|\mathcal{D}^{(\tau)}|$ is a positive even integer. from the original update sequence, and then we apply another reset. The total number of updates between the last two resets is thus $\frac{3}{2}|\mathcal{D}^{(\tau)}|$. By Lemma 7.1, these $\frac{3}{2}|\mathcal{D}^{(\tau)}|$ updates together take time $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau)}|^{\mathsf{w}(\widehat{Q})})$. Moreover, note that between the two resets, the database size cannot change by more than a factor of 2. This is because we only have $\frac{1}{2}|\mathcal{D}^{(\tau)}|$ updates from the original stream that are applied to $\mathcal{D}^{(\tau)}$. In the two extreme cases, all of these updates are inserts or all of them are deletes.
@@
Eq. (63) holds because between $\tau_{i}$ and $\tau_{i+1}$, we are making $\frac{1}{2}|\mathcal{D}^{(\tau_{i})}|$ updates to $\mathcal{D}^{(\tau_{i})}$, which in the two extreme cases are either all inserts or all deletes. As noted before, from the beginning of the $\tau_{i}$-reset until the beginning of the $\tau_{i+1}$-reset, we process $\frac{3}{2}|\mathcal{D}^{(\tau_{i})}|$ updates, which together take time $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau_{i})}|^{\mathsf{w}(\widehat{Q})})$, thanks to Lemma 7.1. To bound the total runtime, we sum up over $i\in[k-1]$: (We don’t actually need to perform the last reset at time $\tau_{k}=N$ since there are no subsequent updates anyhow.)
@@ E.4. Lower Bound for $\text{IVM}^{\pm}_{\delta}[Q]$
The goal of this section is to prove Theorem 4.5, which gives a lower bound on the $\text{IVM}^{\pm}_{\delta}$ update time of a query $Q$ in terms of a lower bound on the static evaluation time for any component $\widehat{Q}_{\bm{\sigma}}$ in its multivariate extension. Specifically, we will prove the following stronger lower bound:
Lemma E.9.
The above lemma immediately implies Theorem 4.5 since the number of single-tuple updates $N$ is larger than the size of the database $|\mathcal{D}|$, and could be unboundedly larger. In order to prove Lemma E.9, we give a reduction from the static evaluation of $\widehat{Q}_{\bm{\sigma}}$ to the $\text{IVM}^{\pm}_{\delta}$ problem for $Q$. Example 7.3 introduces the main idea behind this reduction.